If a company has $300 million of debt and $600 million of equity in its capital structure, what are its debt/total capitalization and debt/equity ratios?
Answer #1

HI,

Here debt = $300 million

equity = $600 million

total capital = debt +equity = 300+600 = $900 million

so debt/total capitalization = 300/900 = 1/3 = 0.33 or 33.33%

debt/equity = 300/600 = 1/2 = 0.5 or 50%

Thanks